E13-7 (L01) (Payroll Tax Entries) Green Day Hardware Companys payroll for November 2017 is summarized below Amount Subject to Payroll Taxes Unemployment Tax Wages Due $120,000 32,000 36,000 $188,000 FICA Federal $40,000 4,000 State Factory Sales Administrative $120,000 32,000 36,000 $188,000 $40,000 4,000 Total $44,000 $44,000 At this point in the year, some employees have already received wages in excess of those to which payroll taxes apply. Assume that the state unemployment taxis 2.5%. The FICA rate is 7.65% on an employees wages to $118,500 and 1.45% in excess of $118,500. Of the $188,000 wages subject to FICA tax, $20,000 of the sales wages is in excess of $118,500. Federal unemployment tax rate is 0.8% after credits. Income tax withheld amounts to $16,000 for factory, $7,000 for sales, and $6,000 for administrative Instructions (a) Prepare a schedule showing the employers total cost of wages for November by function. (Round all computations to nearest dollar.) (b) Prepare the journal entries to record the factory, sales, and administrative payrolls including the employers payroll taxes
Answer #1

Answer

a ) prepare a schedule showing the employer total cost of wages for november by function

factory sales administrative total
salaries and wages 120000 32000 36000 188000
FICA taxes 9180 1208 2754 13142
FUTA taxes 320 32 0 352
SUTA taxes 1000 1000 0 2000
total cost 130500 34240 38754 203494

computation of taxes

factory working
salaries and wages 120000
FICA taxes 9180 7.65% * 120000
FUTA taxes 320 0.8% * 40000
SUTA taxes 1000

2.5% * 40000

total cost 130500
sales
salaries and wages 32000
FICA taxes 1208
FUTA taxes 32 0.8% * 4000
SUTA taxes 100 2.5 % * 4000
total cost 33340

12000 * 7.65 % = 918 , 20000 * 1.45% = 290 , 918 + 290 = 1208

administrative
salaries and wages 36000
FICA taxes 2754 7.65 % * 36000
FUTA taxes 0
SUTA taxes 0
total cost 38754